Hello, can you please explain me the marked parts I know how I got them not sure why I should do them? Like for 2017 sick leave, I thought it was accumulated and for 2017 vacation time it wasn't accumulated and why are we using 2016 projected but 2017 original wage???

*Exercise 13-4 Headland Company began operations on January 2, 2016. It employs 11 individuals who work 8-hour days and are paid hourly. Each employee earns 12 paid vacation days and 7 paid sick days annually Vacation days may be taken after January 15 of the year following the year in which they are earned. Sick days may be taken as soon as they are earned; unused sick days accumulate. Additional information is as follows. Actual Hourly Wage Rate 2016 $12 Vacation Days Used by Each Employee 2016 Sick Days Used by Each Employee 2017 2017 2016 2017 $13 Headland Company has chosen not to accrue paid sick leave until used, and has chosen to accrue vacation time at expected future rates of pay without discounting. The company used the following projected rates to accrue vacation time. Year in Which Vacation Projected Future Pay Rates Time Was Earned 2016 2017 Used to Accrue Vacation Pay $12.58 13.57Prepare journal entries to record transactions related to compensated absences during 2016 and 2017. (If no entry is required, select No Entry for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Round answers to 0 decimal places, e.g. 5,125.) Date Account Titles and Explanation 2016 Salaries and Wages Ex Debit Credit 3284.48 3284.48 laries and Wages Payabl (To accrue the expense and lability for vacations) alaries and Wages Ex 280 28 (To record sick leave paid) No Entry b [114329.921 To record vacation time paid) 2017 Salaries and Wages Ex も alaries and Wages Payabl 4329.92 accrue the alaries and Wages Ex (To record sick leave paid) 0050 만) laries and Wages Ex alaries and Wages Payabl 2177.44 To record vacation time paid) Your answer is correct. Compute the amounts of any liability for compensated absences that should be reported on the balance sheet at December 31, 2016 and 2017. (Round answers to O decimal places, eg. 5,125.) 2016 2017 Accrued liability L32841 15,437
Answer #1

fIRST OF ALL READ THE LAST HALF OF THE QUESTION CAREFULLY IS SPECIFICALY MENTIONS THAT WE HAVE TO ACCRUE VACATION TIME AT FUTURE RATES SO

THAT EXPENSE WE CALCULATE

11(EMPLOYEES) * $13.57 * 8 HOURS A DAY * 12 VACATION =14329.92

FOR SICK PAY WE ARE ACCRUING WHEN THE SICK LEAVE IS ACTUALLY TAKEN SO EXPENSE WILL BE BOOKED 5 DAYS IN 2016 AND 6 DAYS IN 2017 AT ORIGINAL RATES OF 12 & 13 RATHER THAN FUTURE RATES

THIS TABLE SHOULD HELP YOU TO UNDERSTAND

employees per hour hours per day days
(1) 11 × 12.58 × 8 × 12 = 13284.48
(2) 11 × 12 × 8 × 5 = 5280
(3) 11 × 13.57 × 8 × 12 = 14329.92
(4) 11 × 13 × 8 × 6 = 6864
(5) 11 × 12.58 × 8 × 11 = 12177.44
(6) 11 × 13 × 8 × 11 = 12584
64519.84

For the second part

to accrue the liability you are not accruing for sick leave because it is accrued when taken so only vacation pay needs adjustment

in 2016 there was no opening balance for vacation pay , so you accrue all the 12 vacation pays for the coming year at projected rate of 2016 (because you can forecast the rate at 2016 and cannot forecast it for 2017 it will happen only in 2017)

so you accrue 12 vacation days * 8 hours *   11 days *12.58 = $13,284

in 2017 again the same exercise begins you will accrue vacation expense for 12 days plus you have 1 day balance of 2016

again lets make a table and understand:

2016 employees per hour hours per day days
(1) 11 × 12.58 × 8 × 12 = 13284.48
2017
(2) 11 × 12.58 × 8 × 1 = 1107.04
11 × 13.57 × 8 × 12 = 14329.92
15436.96